OpenStudy (anonymous):

Solve the following IVP: y'' -4y' +9y=0 y(0)=0 y'(0)=-8 as you can see it is 2nd order homogeneous differential equation any help please?

OpenStudy (turingtest):

what is the characteristic polynomial?

OpenStudy (anonymous):

r^2-4r+9=0 it is complex root

OpenStudy (turingtest):

and what are those roots?

OpenStudy (anonymous):

2+root5 i 2- root5 i

OpenStudy (anonymous):

characteristic equation is \[D^{2}-4D+9=0\] ITS SOLUTION IS D=\[(4+\sqrt{20}i)\div2\] & \[(4-\sqrt{20}i)\div2\] so the solution is

OpenStudy (turingtest):

the form is\[y(t)=c_1e^{\lambda t}\cos(\mu t)+c_2\sin(\mu t)\]for complex roots\[\lambda\pm\mu i\]so just plug in the corresponding variables what do you then get for the general solution?

OpenStudy (turingtest):

typo, the general solution is of the form\[y(t)=c_1e^{\lambda t}\cos(\mu t)+c_2e^{\lambda t}\sin(\mu t)\]

OpenStudy (anonymous):

ok thanks everyone got it, i just was confused in something now it is all clear thanks

OpenStudy (turingtest):

I'm glad :) welcome!